firm produces uses negotiated transfer prices for its two departments: C and S. Department C transfers its output to Department S. Department S can purchase this output through the external market for $150 per unit.

Department C can sell its output to the external market for $90 per unit. It costs Department C $50 of variable costs per unit. It also costs Department C $20 per cubic foot to ship the output to buyers on the external market. The Department C has $10 of fixed costs per unit.

Department C has 4,000 units of excess capacity. Department S needs 5,000 units.

What is the minimum price that would be negotiated between these departments (round to the nearest cent if necessary)?

a.

$54.00 per unit

b.

$50.00 per unit

c.

$70.00 per unit

d.

$150.00 per unit
Answer #1

Department C has 4,000 units of excess capacity. These 4,000 can be supplied to Department S at the variable cost of $50 per unit.

Hence, transfer price of 4,000 units = 4,000 x 50

= $200,000 (i)

Department S needs 5,000 units.

Remaining 1,000 units would be supplied by Department C recovering its total cost and profit margin.

Profit per unit (If units are sold to external buyers) = Selling price per unit - Total cost per unit

= 90 - (50 + 10 + 20)

= 90 - 80

= $10

Shipment cost of $20 per unit would not be incurred if units are supplied to Department S.

Hence, transfer price of remaining 1,000 units = Variable cost (50) + Fixed cost (10) + Profit margin (10)

= $70 per unit

Hence, transfer price of 1,000 units = 1,000 x 70

= $70,000 (ii)

Total transfer price of 5,000 units = 200,000 + 70,000 [(i) + (ii)]

= $270,000

Hence, transfer price per unit = 270,000/5,000

= $54

Correct option is (a)

Kindly comment if you need further assistance. Thanks